OpenStudy (anonymous):

Is (2, 0) a solution of 2x - y = 4?

hartnn (hartnn):

simply substitute x=2 and y=0,and see whether u get 2x-y as 4

OpenStudy (anonymous):

yes because when u plug in (2,0) into the equation it is a true statement

OpenStudy (anonymous):

2(2)-0=4 4-0=4

OpenStudy (anonymous):

plug it into the x and y values\[2(2)-0=4\]simplify\[4-0=4\]\[4=4?\]^yes ....and so (2,0) is a valid solution :)